# Algorithms with Calibrated Machine Learning Predictions

### v2

@@ C.1 Ski-Rental: CitiBike
(b) Approx end lat.
(c) Approx end long. and lat.
(a) No info about ride end station
(b) Approx end lat.
(c) Approx end long. and lat.
@@
(b) Approx end latitude
(c) Approx end long. and lat.
@@
To calibrate an out-of-the box model, we tested histogram calibration (Zadrozny and Elkan, 2001), binned calibration (Gupta and Ramdas, 2021), and Platt scaling (Platt et al., 1999). While results from histogram and bin calibration were similar, Platt scaling often produced calibrated probabilities within a very small interval. Though it is implemented in our code, we did not use it. A key intervention we make for calibration is to calibrate according to balanced classes in the validation set when the label distribution is highly skewed. This approach ensures that probabilities are not artificially skewed due to class imbalance.

### v3

@@ C.1 Ski-Rental: CitiBike
(a) No info about ride end station
(b) Partial info (approx end latitude)
(c) Rich info (approx end long. and lat.)
@@
(b) Partial info (approx end latitude)
(c) Rich info (approx end long. and lat.)
@@
(b) Partial info (approx end latitude)
(c) Rich info (approx end long. and lat.)
@@
To calibrate an out-of-the box model, we tested histogram calibration (Zadrozny & Elkan, 2001), binned calibration (Gupta & Ramdas, 2021), and Platt scaling (Platt et al., 1999). While results from histogram and bin calibration were similar, Platt scaling often produced calibrated probabilities within a very small interval. Though it is implemented in our code, we did not use it. A key intervention we make for calibration is to calibrate according to balanced classes in the validation set when the label distribution is highly skewed. This approach ensures that probabilities are not artificially skewed due to class imbalance.